OpenStudy (anonymous):

For the function f(x) = (x − 2)2 + 4, identify the vertex, domain, and range. a. The vertex is (–2, 4), the domain is all real numbers, and the range is y ≥ 4. b. The vertex is (–2, 4), the domain is all real numbers, and the range is y ≤ 4. c. The vertex is (2, 4), the domain is all real numbers, and the range is y ≤ 4. d.The vertex is (2, 4), the domain is all real numbers, and the range is y ≥ 4.

OpenStudy (anonymous):

@ali2x2

OpenStudy (michele_laino):

we can rewrite your parabola as below: \[\Large y - 4 = {\left( {x - 2} \right)^2}\]

OpenStudy (michele_laino):

now, if we make this change of variable: \[\Large \left\{ \begin{gathered} Y = y - 4 \hfill \\ X = x - 2 \hfill \\ \end{gathered} \right.\] where X, and Y are the new variables, then your parabola can be rewritten as below: \[\Large Y = {X^2}\]

OpenStudy (michele_laino):

the graph of that parabola is: